Theresa Morgan, a chemist, has a 25% hydrochloric acid solution and a 50% hydrochloric acid solution. How many liters of each should she mix to get 387.5 liters of a hydrochloric acid solution with a 40% acid concentration?



liters of 25% hydrochloric acid solution .
liters of 50% hydrochloric acid solution .

Answer by addingup(3677)

0.25x + 0.50(387.5 - x) = 0.40(387.5)
0.25x + 193.75 - 0.50x = 155
-0.25x = -38.75
x = 155 liters. She needs this number of liters at 25% and 387.5 - 155 = 232.5 liters at 50%
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
232.5(0.5) + 155(0.25) = 155 Correct